Photoelectrons from a material with a binding energy of 2.71 eV are ejected by 450 nm photons. Once ejected, how long (in s) does it take these electrons to travel 2.60 cm to a detection device?
Answer #1

using photoelectric emission

h c / L = phi + kE

0.5* 9.11* 10^-31* v^2 = 6.626* 10^-34* 3* 10^8 / (450* 10^-9) - 2.71* 10^-19* 1.6

v = 1.336* 10^5 m/s

time it takes to cover 2.6 cm

t = 0.026 / 1.336* 10^5

t = 1.946* 10^-7 s

======

comment before rate in case any doubt, will reply for sure.. goodluck